# Hyperbolic recurrent neural network as the first type of non-Euclidean neural quantum state ansatz

### v2

@@ Abstract
In this work, we introduce the first type of non-Euclidean neural quantum state (NQS) ansatz, in the form of the hyperbolic GRU (a variant of recurrent neural networks (RNNs)), to be used in the Variational Monte Carlo method of approximating the ground state wavefunction for quantum many-body systems. In particular, we examine the performances of NQS ansatzes constructed from both conventional or Euclidean RNN/GRU and from hyperbolic GRU in the prototypical settings of the one- and two-dimensional transverse field Ising models (TFIM) and the one-dimensional Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ systems with varying values of $(J_{1},J_{2})$ and $(J_{1},J_{2},J_{3})$. By virtue of the fact that, for all of the experiments performed in this work, hyperbolic GRU can yield performances comparable to or better than Euclidean RNNs, which have been extensively studied in these settings in the literature, our work is a proof-of-concept for the viability of hyperbolic GRU as the first type of non-Euclidean NQS ansatz for quantum many-body systems. Furthermore, in settings where the Hamiltonian displays a clear hierarchical interaction structure, such as the 1D Heisenberg $J_{1}J_{2}$ $J_{1}J_{2}J_{3}$ systems with the first, second and even third nearest neighbor interactions, our results show that hyperbolic GRU definitively outperforms its Euclidean version in almost all instances. Additionally, in the 2D TFIM setting, 1D hyperbolic GRU NQS also outperform 1D Euclidean GRU NQS, due to the presence of a hierarchical interaction structure comprising the first and $N^{th}$ nearest neighbor interactions, introduced by the rearrangement of the 1D spin chain to mimic the 2D lattice. The fact that these results are reminiscent of the established ones from natural language processing where hyperbolic GRU almost always outperforms Euclidean RNNs when the training data exhibit a tree-like or hierarchical structure leads us to hypothesize that hyperbolic GRU NQS ansatz would likely outperform Euclidean RNN/GRU NQS ansatz in quantum spin systems that involve different degrees of nearest neighbor interactions. Finally, with this work, we hope to initiate future studies of other types of non-Euclidean neural quantum states beyond hyperbolic GRU.
@@ 1 Introduction
This paper is organized as follows. In section 2, we recall the details of the Variational Monte Carlo (VMC) method to approximate the ground state energy of a quantum system. In section 3, we summarize the mathematical details of the Poincaré model of hyperbolic space in 3.1 and describe the mathematical constructions of hyperbolic RNN (in 3.2) and of hyperbolic GRU (in 3.3) from their respective Euclidean versions. In section 4, we describe the construction of the neural quantum state trial wavefunction using RNNs, with the real case described in 4.1 and complex case described in 4.2. In section 5, we present the main results obtained by using hyperbolic and Euclidean RNNs as variational ansatzes in the settings of 1D TFIM (section 5.1), 2D TFIM (section 5.2), 1D Heisenberg $J_{1}J_{2}$ model (section 5.3) and 1D Heisenberg $J_{1}J_{2}J_{3}$ model (section 5.4). Section 6 closes the paper with some concluding remarks. In the appendix, we include the convergence curves for the mean energy and variance of all VMC experiments obtained during the training process.
@@ 3.1 Poincaré ball model of hyperbolic space
    $x\oplus_{c}y\equiv\frac{(1+2c\langle x,y\rangle+c||y||^{2})x+(1-c||x||^{2})y}{1+2c\langle x,y\rangle+c^{2}||x||^{2}||y||^{2}}\,.$ (10)
@@
    $\exp_{x}^{c}(v)$ $=$ $x\oplus_{c}\left[\tanh\left(\sqrt{c}\frac{\lambda_{x}^{c}||v||}{2}\right)\frac{v}{\sqrt{c}||v||}\right]$ (11)
    $\log_{x}^{c}(y)$ $=$ $\frac{2}{\sqrt{c}{\lambda}_{x}^{c}}\tanh^{-1}\left(\sqrt{c}||-x\oplus_{c}y||\right)\frac{-x\oplus_{c}y}{||-x\oplus_{c}y||}$ (12)
    where ${\lambda}_{x}^{c}=2/(1-c||x||^{2})$. Note that $\log_{x}^{c}$ and $\exp_{x}^{c}$ as defined in the above equations satisfies $\log_{x}^{c}(\exp_{x}^{c}(v))=v$. When $x=\mathbf{0}$, the above maps take more compact forms
    $\exp_{\mathbf{0}}^{c}(v)$ $=$ $\tanh\left(\sqrt{c}||v||\right)\frac{v}{\sqrt{c}||v||}\,\,\qquad\left(T_{\mathbf{0}_{M}}\mathbb{D}^{M}_{c}\rightarrow\mathbb{D}^{M}_{c}\right)$ (13)
    $\log_{\mathbf{0}}^{c}(y)$ $=$ $\tanh^{-1}\left(\sqrt{c}||y||\right)\frac{y}{\sqrt{c}||y||}\qquad\left(\mathbb{D}^{N}_{c}\rightarrow T_{\mathbf{0}_{N}}\mathbb{D}^{N}_{c}\right)\,.$ (14)
@@
  • The Mobius scalar multiplication of $x\in\mathbb{D}^{N}_{c}\backslash\{\mathbf{0}\}$ by $r\in\mathbb{R}$, denoted by $\otimes_{c}$, is defined as
    $r\otimes_{c}x=\frac{1}{\sqrt{c}}\tanh\left(r\tanh^{-1}(\sqrt{c}||x||)\right)\frac{x}{||x||}$ (17)
@@ 3.2 From Euclidean RNN to Hyperbolic RNN
Given an input vector $\vec{x}_{i}\in\mathbb{R}^{d_{x}}$ at step $i^{th}$ of size $d_{x}$, the defining equation for the Euclidean or conventional version of RNN is one that relates the hidden state vector $h_{i}\in\mathbb{R}^{d_{h}}$ of size $d_{h}$ to the input $x_{i}$ and the same hidden state vector at the previous time step $(i-1)^{\text{th}}$
[…]
@@
[…]
@@
    $W\otimes_{c}(x)=\frac{1}{\sqrt{c}}\tanh\left(\frac{||Wx||}{||x||}\tanh^{-1}\left(\sqrt{c}||x||\right)\right)\frac{Wx}{||Wx||}\,.$ (22)
@@ 3.4 Optimizing hyperbolic parameters
As a final remark in this section, we note that due to the many more mathematical operations involved in defining hyperbolic RNNs compared to their Euclidean counterparts, a drawback of these networks is that they are much more computationally intensive and take much longer to train than Euclidean RNNs. In this work, we only use the GRU variant of hyperbolic RNNs, and not the RNN variant (which tend to underperform GRU) [^6] Throughout this work, however, sometimes we might use the term hyperbolic RNN interchangeably with hyperbolic GRU. This is the case, for example, in the term ‘RNN-based’ where the RNN here can refer to either RNN/GRU (for Euclidean networks) or just GRU (for hyperbolic networks)..
@@ 4.1 Real NQS wavefunction
where $P(\mathbf{\sigma})$ is the output generated by the RNN-based neural network [^8] Applying the RNN quantum state Eq.(36) to the variational energy Eq.(6), we have $E=\sum_{|\vec{\sigma}\rangle}\left(\frac{P(\vec{\sigma})}{\sum_{|\vec{\sigma}\rangle}P(\vec{\sigma})}\right)E_{\text{loc}}(\vec{\sigma})$ (37) where the local energy of a generated sample $\vec{\sigma}$ is $E_{\text{loc}}(\vec{\sigma})=\sum_{|\vec{\sigma}^{\prime}\rangle}\langle\vec{\sigma}|H|\vec{\sigma}^{\prime}\rangle\frac{\sqrt{P(\vec{\sigma}^{\prime})}}{\sqrt{P(\vec{\sigma})}}\,.$ (38) . $P(\sigma)$ is the probability of a particular configuration $|\vec{\sigma}\rangle$, and it is the product of the conditional probabilities $P({\sigma}_{n}|{\sigma}_{1}\ldots s_{n-1})$.
@@
    $\vec{y}_{i}$ is a vector of length $d_{v}$ whose components sum up to 1. The probability $P({\sigma}_{i}|{\sigma}_{1}\ldots s_{i-1})$ is given by
    $P({\sigma}_{i}|{\sigma}_{1}\ldots s_{i-1})=\vec{y}_{i}\,.\,\vec{\sigma}_{i}.$ (43)
@@
Figure 2: The process of generating the length-$N$ samples $\vec{\sigma}=\left({\sigma}_{1},\ldots,{\sigma}_{N}\right)$ using an RNN-based neural network. Each entry ${\sigma}_{i}$ of $\vec{\sigma}$ is sampled from the output $y_{i}$ (generated by the RNN-Dense network using as input the previous entry ${\sigma}_{i-1}$ and the previous RNN hidden state $h_{i-1}$) and then one-hot encoded to be used as input, together with the previous hidden state $h_{i-1}$, for the generation of the next entry ${\sigma}_{i+1}$. For $i=0$, the input $\vec{\sigma}_{0}$ and first hidden state $h_{0}$ are initialized to zero. Figure adapted from [9].
@@ 4.2 Complex NQS wavefunction
Figure 3: Schematic of the calculation of the RNN wavefunction $\Psi(\vec{\sigma})=\sum_{\vec{\sigma}}\exp(i\phi(\vec{\sigma})\sqrt{P(\vec{\sigma})}|\vec{\sigma}\rangle$ where the amplitude $P(\vec{\sigma})=P({\sigma}_{1})P({\sigma}_{2}|{\sigma}_{1})\ldots P(s_{N}|{\sigma}_{N-1})$ and the phase $\phi(\vec{\sigma})=\sum_{i=1}^{N}\phi(\vec{\sigma}_{i})$. For ease of illustration, $N$ is chosen to be 3. (S) and (SS), correspondingly, denote the dense layer with the Softmax and Softsign activation function. Figure adapted from [9].
@@ 5 VMC with hyperbolic RNN-based NQS
These systems are the most basic testing grounds for different types of NQS ansatzes, and they have been used routinely in the literature on NQS. As benchmark against which to evaluate the performances of all NQS ansatzes, we will use DMRG (Density Matrix Renormalization Group). With 1D systems, DMRG results are considered to be exact, while in 2D systems, this is no longer the case, but this is still a reliable benchmark to use since DMRG provides state-of-the-art computational results for various condensed matter systems.
@@
As the performances of Euclidean RNN-based NQS have been established in the work [9] for the 1D TFIM, 2D TFIM and 1D $J_{1}J_{2}$ systems, we do not aim to replicate their results but instead aiming to establish the performances of hyperbolic GRU-based NQS using the performances of Euclidean RNN/GRU-based NQS as benchmarks in settings of smaller scales, due to the fact that hyperbolic networks are more computationally intensive and take much longer to train than their Euclidean counterparts. Specifically, for the 1D TFIM , the work [9] explored settings up to $N=1000$ spins, while we restrict the number of spins in our settings up to $N=100$. For the 2D TFIM, [9] studied a square lattice of size $(N_{x},N_{y})=(12,12)$ while we restrict the lattice size to $(N_{x},N_{y})=(9,9)$. For the $J_{1}J_{2}$ systems, [9] used a setting involving $N=100$ spins while we restrict the number of spins in our setting to $N=50$. The 1D $J_{1}J_{2}J_{3}$ model was not studied in [9].

### v3

@@ Abstract
In this work, we introduce the first type of non-Euclidean neural quantum state (NQS) ansatz, in the form of the hyperbolic GRU (a variant of recurrent neural networks (RNNs)), to be used in the Variational Monte Carlo method of approximating the ground state wavefunction for quantum many-body systems. In particular, we examine the performances of NQS ansatzes constructed from both conventional or Euclidean RNN/GRU and from hyperbolic GRU in the prototypical settings of the one- and two-dimensional transverse field Ising models (TFIM) and the one-dimensional Heisenberg $J_{1}J_{2}$ and $J_{1}J_{2}J_{3}$ systems with varying values of $(J_{1},J_{2})$ and $(J_{1},J_{2},J_{3})$. By virtue of the fact that, for all of the experiments performed in this work, hyperbolic GRU can yield performances comparable to or better than Euclidean RNNs, which have been extensively studied in these settings in the literature, our work is a proof-of-concept for the viability of hyperbolic GRU as the first type of non-Euclidean NQS ansatz for quantum many-body systems. Furthermore, in settings where the Hamiltonian displays a clear hierarchical interaction structure, such as the 1D Heisenberg $J_{1}J_{2}$ & $J_{1}J_{2}J_{3}$ systems with the first, second and even third nearest neighbor interactions, our results show that hyperbolic GRU definitively outperforms its Euclidean version in almost all instances. Additionally, in the 2D TFIM setting, 1D hyperbolic GRU NQS also outperforms 1D Euclidean GRU NQS, due to the presence of a hierarchical interaction structure comprising the first and $N^{th}$ nearest neighbor interactions, introduced by the rearrangement of the 1D spin chain to mimic the 2D lattice. The fact that these results are reminiscent of the established ones from natural language processing where hyperbolic RNNs almost always outperform Euclidean RNNs when the training data exhibit a tree-like or hierarchical structure leads us to hypothesize that hyperbolic GRU NQS ansatz would likely outperform its Euclidean counterpart in quantum spin systems that involve different degrees of nearest neighbor interactions. Finally, with this work, we hope to initiate future studies of other types of non-Euclidean neural quantum states beyond hyperbolic GRU.
@@ 1 Introduction
This paper is organized as follows. In section 2, we recall the details of the Variational Monte Carlo (VMC) method to approximate the ground state energy of a quantum system. In section 3, we summarize the mathematical details of the Poincaré model of hyperbolic space in 3.1 and describe the mathematical constructions of hyperbolic RNN (in 3.2) and of hyperbolic GRU (in 3.3) from their respective Euclidean versions. In section 4, we describe the construction of the neural quantum state trial wavefunction using RNNs, with the real case described in 4.1 and complex case described in 4.2. In section 5, we present the main results obtained by using hyperbolic and Euclidean RNNs as variational ansatzes in the settings of 1D TFIM (section 5.1), 2D TFIM (section 5.2), 1D Heisenberg $J_{1}J_{2}$ model (section 5.3) and 1D Heisenberg $J_{1}J_{2}J_{3}$ model (section 5.4). Section 6 closes the paper with some concluding remarks. In the appendix, we include the convergence curves for the mean energy, obtained during the training process, of all VMC experiments.
@@ 3.1 Poincaré ball model of hyperbolic space
[…]
@@
[…]
@@
[…]
@@ 3.2 From Euclidean RNN to Hyperbolic RNN
Given an input vector $\vec{x}_{i}\in\mathbb{R}^{d_{x}}$ at step $i$ of size $d_{x}$, the defining equation for the Euclidean or conventional version of RNN is one that relates the hidden state vector $h_{i}\in\mathbb{R}^{d_{h}}$ of size $d_{h}$ to the input $x_{i}$ and the same hidden state vector at the previous time step (i1)(i-1)
$\text{RNN}:$ $\vec{h}_{i}=f(W_{h}\vec{h}_{i-1}+U_{h}\vec{x}_{i}+\vec{b}_{h})\,$, (19)
@@
$\text{Hyperbolic RNN}:$ $\vec{h}_{i}=f^{\otimes_{c}}\left[(W_{h}\otimes_{c}\vec{h}_{i-1})\oplus_{c}(U_{h}\otimes_{c}\vec{x}_{i})\oplus_{c}\frac{}{}\vec{b}_{h}\right]$ (20)
@@
[…]
@@ 3.4 Optimizing hyperbolic parameters
As a final remark in this section, we note that due to the many more mathematical operations involved in defining hyperbolic RNNs compared to their Euclidean counterparts, a drawback of these networks is that they are much more computationally intensive and take much longer to train than Euclidean RNNs. In this work, we only use the GRU variant of hyperbolic RNNs, and not the RNN variant (which tends to underperform GRU) [^6] Throughout this work, however, sometimes we might use the term hyperbolic RNN interchangeably with hyperbolic GRU. This is the case, for example, in the term ‘RNN-based’ where the RNN here can refer to either RNN/GRU (for Euclidean networks) or just GRU (for hyperbolic networks)..
@@ 4.1 Real NQS wavefunction
where $P(\mathbf{\sigma})$ is the output generated by the RNN-based neural network [^8] Applying the RNN quantum state Eq.(36) to the variational energy Eq.(6), we have $E=\sum_{|\vec{\sigma}\rangle}\left(\frac{P(\vec{\sigma})}{\sum_{|\vec{\sigma}\rangle}P(\vec{\sigma})}\right)E_{\text{loc}}(\vec{\sigma})$ (37) where the local energy of a generated sample $\vec{\sigma}$ is $E_{\text{loc}}(\vec{\sigma})=\sum_{|\vec{\sigma}^{\prime}\rangle}\langle\vec{\sigma}|H|\vec{\sigma}^{\prime}\rangle\frac{\sqrt{P(\vec{\sigma}^{\prime})}}{\sqrt{P(\vec{\sigma})}}\,.$ (38) . $P(\sigma)$ is the probability of a particular configuration $|\vec{\sigma}\rangle$, and it is the product of the conditional probabilities $P({\sigma}_{n}|{\sigma}_{1}\ldots{\sigma}_{n-1})$.
@@
    $\vec{y}_{i}$ is a vector of length $d_{v}$ whose components sum up to 1. The probability $P({\sigma}_{i}|{\sigma}_{1}\ldots{\sigma}_{i-1})$ is given by
    $P({\sigma}_{i}|{\sigma}_{1}\ldots{\sigma}_{i-1})=\vec{y}_{i}\,.\,\vec{\sigma}_{i}.$ (43)
@@
Figure 2: The process of generating the length-NN samples $\vec{\sigma}=\left({\sigma}_{1},\ldots,{\sigma}_{N}\right)$ using an RNN-based neural network. Each entry ${\sigma}_{i}$ of $\vec{\sigma}$ is sampled from the output $y_{i}$ (generated by the RNN-Dense network using as input the previous entry ${\sigma}_{i-1}$ and the previous RNN hidden state $h_{i-1}$) and then one-hot encoded to be used as input, together with the previous hidden state $h_{i-1}$, for the generation of the next entry ${\sigma}_{i+1}$. For $i=0$, the input $\vec{\sigma}_{0}$ and first hidden state $h_{0}$ are initialized to zero. Figure adapted from [9].
@@ 4.2 Complex NQS wavefunction
Figure 3: Schematic of the calculation of the RNN wavefunction $\Psi(\vec{\sigma})=\sum_{\vec{\sigma}}\exp(i\phi(\vec{\sigma})\sqrt{P(\vec{\sigma})}|\vec{\sigma}\rangle$ where the amplitude $P(\vec{\sigma})=P({\sigma}_{1})P({\sigma}_{2}|{\sigma}_{1})\ldots P(s_{N}|{\sigma}_{N-1})$ and the phase $\phi(\vec{\sigma})=\sum_{i=1}^{N}\phi(\vec{\sigma}_{i})$. For ease of illustration, NN is chosen to be 3. (S) and ($S$), correspondingly, denote the dense layer with the Softmax and Softsign activation function. Figure adapted from [9].
In the experiments involving the complex NQS wavefunction in this work, the Marshall sign [47] is used (since it leads to some improvements in the results, as reported by [9]). The wavefunction with the Marshall sign, $\Psi_{M}(\vec{\sigma})$, can be written as
$\Psi_{M}(\vec{\sigma})=(-1)^{M_{A}(\vec{\sigma})}\Psi(\vec{\sigma})$ (50)
where $M_{A}(\vec{\sigma})=\sum_{i\in A}{\sigma}_{i}$, with $A$ comprising the sites belonging to the sublattice of all even or all odd sites in the lattice. The factor $(-1)^{M_{A}(\vec{\sigma})}$ is the Marshall sign of the wavefunction, and $\Psi({\sigma})$ is the NQS wavefunction before the Marshall sign was applied (for complex NQS, this comprises both the amplitude and phase parts). Eq.(50) is exact in unfrustrated spin systems when only neareast neighbor interactions $J_{1}$ (among the spin sites) are nonzero. When next-nearest (or higher degrees of) neighbor interactions ($J_{2}$ or $J_{i\geq 2}$) are nonzero, Eq.(50) is no longer exact but still holds approximately [7].
@@ 5 VMC with hyperbolic RNN-based NQS
These systems are the most basic testing grounds for different types of NQS ansatzes, and they have been used routinely in the literature. As a benchmark against which to evaluate the performances of all NQS ansatzes, we will use DMRG (Density Matrix Renormalization Group). With 1D systems, DMRG results are considered to be exact, while in 2D systems, this is no longer the case, but this is still a reliable benchmark to use since DMRG provides state-of-the-art computational results for various condensed matter systems.
@@
As the performances of Euclidean RNN-based NQS have been established in the work [9] for the 1D TFIM, 2D TFIM and 1D $J_{1}J_{2}$ systems, we do not aim to replicate their results but instead to establish the performances of hyperbolic GRU-based NQS using the performances of Euclidean RNN/GRU-based NQS as benchmarks in settings of smaller scales, due to the fact that hyperbolic networks are more computationally intensive and take much longer to train than their Euclidean counterparts. Specifically, for the 1D TFIM , the work [9] explored settings up to $N=1000$ spins, while we restrict the number of spins in our settings up to $N=100$. For the 2D TFIM, [9] studied a square lattice of size $(N_{x},N_{y})=(12,12)$ while we restrict the lattice size to $(N_{x},N_{y})=(9,9)$. For the $J_{1}J_{2}$ systems, [9] used a setting involving $N=100$ spins while we restrict the number of spins in our setting to $N=50$. The 1D $J_{1}J_{2}J_{3}$ model was not studied in [9].
Before moving further, several important remarks are in order. These concern the different aspects of the experiments that remain unchanged throughout this work (unless otherwise explicitly stated).
  • During the training process of all neural networks in this work, we fixed the number of generated spin configuration samples to be 50. At each iteration of the training process, a neural network ansatz generates 50 samples (spin configurations) whose mean energy is used to approximate the true ground state energy of the Hamiltonian system under study. Convergence is reached when there is no noticeable change in the mean energy and the energy variance of all the samples is smaller than a certain tolerance threshold [^10] This specific values of the tolerance threshold vary for different VMC experiments, and can range from as low as 0.05 to 1.0 for the 1D TFIM case. In other settings such as 2D TFIM, 1D $J_{1}J_{2}$ and 1D $J_{1}J_{2}J_{3}$ models, the variance tolerance can be much higher.. This choice of 50 samples (during training) was made based on the known fact that RNN-based, autoregressive NQS do not require too many samples to achieve convergence since an autoregressive architecture allows for an exact and independent sampling process by forward-generating spins from conditional distributions (with no Markov chain Monte Carlo sampling required) [5], [9]. Furthermore, we also observed during our experiments that increasing the number of samples generated to 100 or 200 did not really improve the results while exponentially increasing the training time [^11] For Euclidean RNN/GRU, we had no problem using 200 samples during the training process, and 200 was in fact the number of samples used in the work [9], but for hyperbolic GRU, this took too long with the limited computing resources that we currently have. Already with 50 samples, the training of hyperbolic GRU took much longer than that of their Euclidean counterparts.. In the inference process, however, when the trained neural networks were loaded to generate new samples to estimate the energy, we used a significantly higher number (in the order of thousands) of samples to estimate the mean energy.
  • For Euclidean ansatzes, we used Adam optimizer with an exponential decay learning rate while for the hyperbolic ansatz, we used RSGD (Riemannian SGD) with the fixed learning rate of ${\alpha}=10^{-2}$ for hyperbolic parameters and the same Adam with exponentially decaying learning rate for Euclidean parameters. In the hyperbolic GRU ansatz, only the biases ($b_{h},b_{r},b_{z})$ are hyperbolic, while the weight matrices are Euclidean. This hyperbolic network configuration is the same as the one used in [30], and is kept unchanged throughout this work.
  • In a typical training process, the weights of the neural network model under training are constantly overwritten after each training epoch, so the final model is the one saved at the last epoch in the training. Theoretically, at convergence when the model reaches a constant value without any further changes, the weights of the model stop changing and the weights at the last epoch are the final weights of the model. In practice, what is observed is that even at convergence when the model has reached the constant energy level, there are still minor fluctuations within a very small range of this energy, espcially with this kind of training involving the generation of samples. To address this issue, in our training process, the saving of the model weights at any particular training epoch is conditioned on the fact that the model has actually improved (the criteria for judging the improvement, programmed in the training loop, include not only a lower mean energy but also a lower variance, subject to a certain variance tolerance threshold) compared to the previous training epoch. As such, when convergence is reached in a practical sense (in the sense that the mean energy has reached a constant level albeit with fluctuations within a very narrow range), our last saved model correponds to the best possible result, i.e. a model that has already converged with the lowest energy and the variance smaller than the tolerance threshold. This means that our final model might not necessarily be the model at the last epoch in the training, but the model saved at some earlier epoch near the end of the training [^12] We note that the practice of saving the best model only after a practical convergence is reached is commonly used in all types of neural network training not restricted to the present context of NQS. In Keras (keras.io/api/callbacks/model_checkpoint/), one can choose to save the best model only (e.g. one that has the lowest validation loss or one that fulfills a certain set of custom defined criteria) as the training progresses using a function built-in in ModelCheckpoint callback by selecting the option save_best_only = True. . In other words, the training loop might continue until the specified number of iterations is reached but unless a better result is obtained, the final model is the model saved at the epoch where the best result was obtained [^13] Our training results are perfectly reproducible, as long as the same random seed parameter is chosen. . For example, with the total number of epochs set at 450, if the convergence is reached after 350 epochs (the energy no longer changes noticeably but varies within a very narrow range of order 0.01 to 0.1), and the best model (with the lowest energy and the variance below the threshold) is the one at epoch 375, training will still carry on until epoch 450 but no new model is saved after epoch 375 unless there is an improved model at the $i^{th}$ epoch where $375<i<450$.